### v1

# Enhanced Generative Recommendation via Content and Collaboration Integration

Abstract.
Generative recommendation has emerged as a promising paradigm aimed at augmenting recommender systems with recent advancements in generative artificial intelligence. This task has been formulated as a sequence-to-sequence generation process, wherein the input sequence encompasses data pertaining to the user’s previously interacted items, and the output sequence denotes the generative identifier for the suggested item. However, existing generative recommendation approaches still encounter challenges in (i) effectively integrating user-item collaborative signals and item content information within a unified generative framework, and (ii) executing an efficient alignment between content information and collaborative signals.
In this paper, we introduce content-based colla-borative generation for recommender systems, denoted as ColaRec. To capture collaborative signals, the generative item identifiers are derived from a pretrained collaborative filtering model, while the user is represented through the aggregation of interacted items’ content. Subsequently, the aggregated textual description of items is fed into a language model to encapsulate content information. This integration enables ColaRec to amalgamate collaborative signals and content information within an end-to-end framework. Regarding the alignment, we propose an item indexing task to facilitate the mapping between the content-based semantic space and the interaction-based collaborative space. Additionally, a contrastive loss is introduced to ensure that items with similar collaborative GIDs possess comparable content representations, thereby enhancing alignment. To validate the efficacy of ColaRec, we conduct experiments on three benchmark datasets. Empirical results substantiate the superior performance of ColaRec.
Recommender System, Generative Recommendation, Collaborative Filtering, Generative Model, Language Model
isbn: ccs: Information systems Recommender systemsccs: Information systems Retrieval models and rankingccs: Information systems Novelty in information retrieval
1. introduction
Recommender systems are widely deployed to discover user interests and provide personalized information services (Hu et al., 2018; Qi et al., 2020; Yuan et al., 2019). Recently, generative models, such as large language models (LLMs) (Zhao et al., 2023; Chang et al., 2023) and diffusion models (Yang et al., 2023; Cao et al., 2022), have gained prominence in advancing artificial intelligence. In such a context, generative recommendation (Wang et al., 2023b; Bao et al., 2023; Si et al., 2023; Rajput et al., 2023; Geng et al., 2022; Ji et al., 2023) has emerged as a new paradigm to enhance recommendation agents with end-to-end generation capability of these models.
Generative Recommendation. Generative recommendation assigns each item with a unique sequence of tokens as the item’s generative identifier, a.k.a., GID, then the recommendation is formulated as a sequence-to-sequence generation task, where the input is historical user-item interactions while the output sequence refers to the GID of the recommended item (Rajput et al., 2023; Si et al., 2023; Hua et al., 2023) [^1] Wang et al. (2023a) proposed another paradigm to directly generate new content, e.g., images, for recommendation. However, their methods are tailored for the generation of virtual content and cannot be used for recommendation of concrete items, such as the item recommendation in an e-commerce platform. In this paper, we target on the recommendation of concrete items..
Figure 1. Comparison between conventional itemIDs and GIDs. GIDs contain more concrete correlations.
Compared with conventional itemIDs with an assigned single random token, the sequential tokens of a GID contain more explicit representation information. As shown in Figure 1, we see correlated GIDs denote correlated items, and thus help the recommender to gain better performance. Generative recommendation provides an end-to-end paradigm to effectively utilize generative models for recommendation. Tay et al. (2022) initially introduced Differentiable Search Indexing (DSI), which marks an early endeavor in employing generative models for information retrieval. DSI has inspired a series of studies on generative recommendation (Rajput et al., 2023; Si et al., 2023; Petrov and Macdonald, 2023). The promising results of these works demonstrate the potential of generative recommendation.
Limitations of Current Approaches. Existing generative recommendation approaches still have the following two limitations, leading to suboptimal performance.
Firstly, existing methods fail to effectively model collaborative signals and item content information in a unified framework. Collaborative signals refer to the knowledge contained in the user-item interactions while item content information refers to the textual description of items, as shown in Figure 2. On one hand, several methods utilize either the item title or hierarchical content embeddings of items to construct GIDs (Li et al., 2023c; Rajput et al., 2023). However, these methods only consider the item content information while the collaborative signals between users and items are overlooked. On the other hand, Si et al. (2023) construct GIDs using item embeddings of a pretrained SASRec (Kang and McAuley, 2018). Although the GID in this work contains the item-item sequential (collaborative) connections, the proposed recommendation framework in their research fails to model the item content information. To achieve satisfying recommendation results, we argue that the generative recommenders should be able to jointly model the item content information and the user-item collaborative signals in a unified generative framework.
Secondly, existing generative recommendation methods cannot effectively align item content information and collaborative signals. Although Hua et al. (2023) proposed to construct the GID using the concatenation of a semantic string learnt from the item content and a sequence of tokens learnt from the item-item co-occurrence matrix, the naive concatenation without a learning process is not effective to perform the alignment, leading to sub-optimal performance. The rich content information can assist the recommendation models in obtaining more informative and comprehensive item representations, thus enhancing the modeling of user-item collaborative interactions, and vice visa. To this end, the alignment between collaborative signals and content information is fundamental for further enhancing generative recommenders. However, we argue that a successful alignment should be achieved through an explicit learning process, e.g.,the mapping between the content-based semantic space and the interaction-based collaborative space.
Figure 2. Illustration of collaborative signals and content information. Collaborative signals refer to the knowledge contained in user-item interactions while content information refers to the textual description of items.
Proposed Methods. In this paper, we propose content-based colla-borative generation for recommender systems (ColaRec), a generative recommendation model which unifies both item content information and user-item collaborative interaction signals in a sequence-to-sequence generation framework. Taking a particular user as an example, the input sequence of ColaRec consists of unordered [^2] In this paper, we target on the general recommendation task other than sequential recommendation. To this end, we use unordered item tuples as the input. tuples with each tuple describing the content information of one interacted item of this user. Then an encoder-decoder framework is used to generate the GID of the target item.
To address the first limitation, we propose two tailored designs to jointly model user-item collaborative signals and item content information. Firstly, we propose to construct GIDs using item representations obtained from a pretrained collaborative filtering model. In this paper, we use LightGCN (He et al., 2020) as the pretrained model. The LightGCN model is trained on the user-item interaction graph and thus the constructed GID can effectively encodes the user-item collaborative signals. Note that the LightGCN model can also be alternated with other models. Secondly, in ColaRec, the user is represented as the content aggregation of tuples with each tuple describing one her/his historically interacted item. This design also keeps inline with the nature of collaborative filtering. The aggregation of content-based tuples is fed to an encoder-decoder based language model to effectively capture the textual content information. In this paper, we use a fine-tuned T5 (Raffel et al., 2020) as the involved language model. We leave the investigation of larger language models as one of our future works.
To address the second limitation, we propose an auxiliary item indexing task which targets on mapping the item side information into the GID of this item through the same encoder-decoder model. More precisely, the item side information contains both the textual content information and a set of users who have interacted with this item. To this end, the indexing task maps both the item content information and user-item interaction signals into the constructed GID, achieving better alignment between item content information and user-item collaborative signals. Besides, we further propose a contrastive loss to ensure that items with similar collaborative GIDs are also similar in the content-based semantic space. Such the contrastive loss also helps to perform alignment between content information and collaborative signals.
To demonstrate the effectiveness of the proposed ColaRec, we conduct extensive experiments on three public accessible benchmark datasets. Experimental results show that the proposed ColaRec outperforms related state-of-the-art baselines.
Main Contributions. Our main contributions are as follows:
  • We propose ColaRec, an end-to-end generative recommender which utilizes an encoder-decoder based language model to jointly model item content information and user-item collaborative signals for recommendation.
  • We propose an auxiliary item indexing task and a contrastive loss to perform better alignment between item content information and user-item collaborative signals to further enhance the performance of generative recommendation.
  • We conduct extensive experiments on three datasets to demonstrate the effectiveness of the proposed ColaRec. Experimental results show superior recommendation performance of ColaRec.
2. related work
In this section, we review related prior literature on collaborative filtering and generative models.
2.1. Collaborative Filtering
Collaborative filtering (CF) is one of the most representative methods to build a recommendation agent. CF believes that a user can be represented as the aggregation of her/his interacted items, and vice visa. The keystone to conduct collaborative filtering is the user-item interaction matrix. Early approaches (Koren et al., 2009; Rendle et al., 2012) are based on matrix factorization (MF) to jointly model the latent space for users and items. Due to the expressiveness of deep neural networks, plenty of research (He et al., 2017; Mao et al., 2021; Zheng et al., 2017; Chen et al., 2018; Gao et al., 2020) has been conducted to enhance CF through deep learning. Besides, since the user-item interaction signals can be naturally encoded into an interaction graph, graph neural networks (GNN) also shed lights in the field of CF. Berg et al. (2017) proposed to use graph convolution for matrix completion. Wang et al. (2019) proposed the NGCF model to use GNN for CF. He et al. (2020) proposed the LightGCN model to simplify GNN for recommender systems, leading to a simpler and linear CF model. LightGCN serves as one of the most popular GNN-based CF approaches due to its effectiveness. Lin et al. (2022) proposed NCL to introduce contrastive learning into graph-based CF. Besides, content information has also been utilized to enhance the CF model. Rendle (2010) proposed the notable factorization machine (FM) to extend MF for categorical contextual features. He and McAuley (2016); Wei et al. (2019, 2020); Wu et al. (2017) proposed to enhance CF with visual or text features. Li et al. (2023b) proposed RecFormer to model long text sequences for recommendation.
Different with existing CF methods, in this paper we focus on the new paradigm of generative recommendation, where the recommended item is generated through the decoder of a sequence-to-sequence model. Generative recommendation provides new potential to utilize advances in the field of generative artificial intelligence to enhance the recommendation task.
2.2. Generative Models
Generative models have become a hot research topic to generate new content, such as images and text, from the collected data. Variational Autoencoders (VAEs) (Kingma and Welling, 2014; Higgins et al., 2016) and Generative Adversarial Networks (GANs) (Goodfellow et al., 2014; Karras et al., 2019; Creswell et al., 2018), are two kinds of representative generative models. VAEs aim to learn a probabilistic mapping from the observed data to a latent space, and then leverage variational inference to approximate the posterior distribution of latent variables. GANs introduce a min-max gaming approach with a generator attempting to generate simulated samples, and the discriminator distinguishing between real and generated data. Recently, Transformers (Vaswani et al., 2017) have shown promise in language generation, leading to notable LLMs like GPTs (Generative Pre-trained Transformers). Besides, diffusion models (Ho et al., 2020; Yang et al., 2023; Cao et al., 2022) have also shown promising results in content generation.
Generative models for document retrieval. Generative retrieval has drawn increasing attention to conduct document retrieval with a sequence-to-sequence framework, where the input sequence is the query and the output sequence is the document identifier. The bloom of LLMs encourages generative retrieval to better understand semantic patterns stored in the retrieval process. Tay et al. (2022) firstly proposed a differentiable search index (DSI) for document retrieval, which assigns each document with a structured semantic meaningful identifier. DSI can be regarded as the early attempt to conduct generative retrieval. Plenty of work has been conducted following this research line. For example, strings including titles (Tang et al., 2023; Cao et al., 2021), n-grams (Bevilacqua et al., 2022; Chen et al., 2023) or URLs (Zhou et al., 2022) are explored to construct the document identifier. Wang et al. (2022a) proposed NCI which produces a hierarchical document identifier to get fine-grained semantic mapping, and a novel decoder to consider the prefixes of document identifiers. GenRet (Sun et al., 2023) learns to tokenize the document into a short discrete representation via a discrete auto-encoding approach, and utilizes progressive training and diverse clustering to assist the model training process.
Figure 3. Overview of ColaRec. ColaRec assigns each item with a GID obtained from a GNN-based CF model. ColaRec consists two tasks. User-item recommendation aims to map the user’s interacted items with textual content into the GID of the recommended item, i.e., $\mathcal{L}_{\text{rec}}$. Item-item indexing targets on the mapping from item side information into the item’s GID, i.e., $\mathcal{L}_{\text{index}}$. Besides, a ranking loss $\mathcal{L}_{\text{bpr}}$ and a contrastive loss $\mathcal{L}_{\text{c}}$ are also introduced. A tuned T5 is used to accomplish these tasks.
Generative models for recommendation. Generative models have also been utilized for recommender systems. Early works focus on utilizing VAEs (Liang et al., 2018; Shenbin et al., 2020; Cai and Cai, 2022) or GANs (He et al., 2018; Guo et al., 2022; Wang et al., 2022b) for recommendation. Recently, diffusion models have also been utilized to perform item recommendation (Wang et al., 2023b; Li et al., 2023a). Besides, LLM-based recommendation has also been proposed (Bao et al., 2023; Zhang et al., 2023; Petrov and Macdonald, 2023). These works explore the ability of LLMs for recommendation through parameter efficient tuning (Bao et al., 2023) or instruction tuning (Zhang et al., 2023). Meanwhile, several works attempt to convert various recommendation tasks into a unified natural language generation task, and train recommenders with multi-task optimization, like P5 (Geng et al., 2022) and M6-Rec (Cui et al., 2022).
Notably, TIGER (Rajput et al., 2023) is a representative generative recommendation method which uses an RQ-VAE (Zeghidour et al., 2022) to construct the GID and then uses encoder-decoder based transformers to generate sequential recommendation. Si et al. (2023) proposed to construct the GID from a pretrained SASRec (Kang and McAuley, 2018) model for sequential recommendation. Hua et al. (2023) further investigated the effect of item identifier construction for generative recommendation.
Despite the arise of generative recommenders, existing methods still suffer from the ineffective infusion of collaborative signals and item content information. How to jointly model and align collaborative signals and content information in a sequence-to-sequence generative framework for recommender systems is still an open research challenge.
3. Notations and Task Formulation
We first introduce the notations used in this paper. Then, we describe the task definition of generative recommendation.
3.1. Notations
[…]
3.2. Generative Recommendation
The task of generative recommendation is given the input describing the information of $\mathcal{I}_{u}^{+}$, generating a list of GIDs as the recommendation result. The GID is generated through an auto-regressive manner. The probability of recommending item $i$ for user $u$ is estimated as:
(1) $p(u,i)=\prod_{t=1}^{l}p(z_{i}^{t}|z_{i}^{1},z_{i}^{2},\cdots,z_{i}^{t-1},\mathcal{I}_{u}^{+}).$
The recommender selects items with the top-$n$ highest $p(u,i)$ as the recommendation list for user $u$.
4. Methodology
In this section, we describe the details of ColaRec. We first provide overview of the proposed ColaRec. Then the construction of items’ GIDs are detailed. After that, the user-item recommendation task and the item-item indexing task are described. Finally, we describe the joint optimization of above tasks.
4.1. Overview of ColaRec
Figure 3 illustrates an overview of the proposed ColaRec. ColaRec constructs the GID using a graph-based CF model, which effectively captures collaborative signals. The training of ColaRec consists of two tasks: the user-item recommendation task and the item-item indexing task. The user-item recommendation aims to map the content information of the user’s historical interacted items into the GID of the recommended item. The item-item indexing task targets on the mapping from the item side information into the item’s GID. Both the two tasks are achieved through a shared encoder-decoder based language model to better capture the textual content information. To this end, the recommendation task unifies both collaborative signals and item content information for better recommendation, while the indexing task performs the alignment between collaborative signals and content information. Note that parameters of the involved language model are also fine-tuned to better adapt the language model for recommendation.
4.2. Generative Identifier Construction
The construction of GIDs plays a crucial role for generative recommendation. Generally speaking, GIDs should satisfy the following expectations for better recommendation: (i) GIDs need to contain knowledge about both collaborative signals and content information; (ii) correlated items (e.g., similar items in content or items interacted by similar users) should have correlated GIDs; (iii) each item should have one unique GID and each GID should correspond to one specific item.
To fulfil above expectations, we utilize a hierarchical clustering approach to construct GIDs from a graph-based CF model. Specifically, we first extract item representations from a pretrained LightGCN model. Then the constrained $K$-means algorithm is called hierarchically based on the item representations. For the $t$-th level clustering with $t\in[1,l-1]$, the number of clusters is set to $K^{l-t}$. The next level clustering is conducted with items in the current cluster as the whole instance set. Regarding the last level of leaf node, we randomly allocate 1 to $K$ to the item. In this way, we establish a $K$-ary tree to organize the item set. Each item corresponds to a leaf node, while the path from the root to the leaf node is the GID of the item. Since LightGCN is trained on the user-item interaction graph, the GID can naturally encodes collaborative signals. Meanwhile, there is a codebook embedding matrix for every position of the GID, which will incorporate the content information in the item indexing task. We give detailed description in section 4.4. To this end, the GID together with the corresponding codebook embedding helps the recommender to model both collaborative signals and content information.
4.3. User-Item Recommendation
Model Inputs. The input sequence of user-item recommendation for user $u$ consists of unordered tuples with each tuple describing the content information of one interacted item of this user. For textual description of item $i$, we adopt the universal data format from (Li et al., 2023b). Specifically, textual description $c_{i}$ of item $i$ is formulated as an item “sentence” that comes from a flatted attribute dictionary consisting of key-value attribute pairs $(k,v)$, i.e., $[k_{1}{:}v_{1},k_{2}{:}v_{2},\dots]$. Besides, we also introduce the item atomic identifier $iad_{i}$ into the content information to further increase model fidelity. To this end, the content tuple of item $i$ is formulated as:
(2) $c_{i}=[iad_{i},k_{1}{:}v_{1},k_{2}{:}v_{2},\dots].$
The key idea of CF is that users’ preferences can be inferred from their interacted items. Therefore, for each user $u$, the input consists of the content aggregation of item tuples that $u$ has interacted with to reinforce the collaborative signals.
Since the training of ColaRec has two tasks, in the user-item recommendation task, we augment the input by adding a textual “user prompt” (denoted as $prompt_{u}$) at the beginning of the input to inform the model that the ongoing task is the recommendation task. Thus, the input for the user-item recommendation task is:
(3) $X_{u}=[prompt_{u},\{c_{i}|i\in\mathcal{I}_{u}^{+}\}].$
Language Models for Recommendation. We employ an encoder-decoder based language model to capture textual content information. Given the model input $X_{u}$, the language model encoder captures the semantic information of $X_{u}$ and returns the hidden state $\text{Encoder}(X_{u})$. After that, given the generated tokens $z^{<t}$ before the $t$-th generation step, the decoder generates the latent representation $\mathbf{d}_{t}\in\mathbb{R}^{m}$ for the $t$-th token of GID. $m$ is the dimension of the latent representation. This process can be formulated as:
(4) $\mathbf{d}_{t}=\text{Decoder}(\text{Encoder}(X_{u}),z^{<t})$
The generation probability at step $t$ is estimated by $\mathbf{d}_{t}$ and the codebook embedding matrix for position $t$, which is formulated as:
(5) $p(z^{t}|z^{<t},X_{u})=\operatorname{softmax}(\mathbf{d}_{t}\cdot\mathbf{E}_{t}^{\top}),$
where $\mathbf{E}_{t}$ is the $t$-th step codebook embedding matrix.
We adopt the cross-entropy loss for model optimization. Specifically, given a $(u,i)$ pair in the training set, the generative loss of recommendation is formulated as:
(6) $\mathcal{L}_{\text{rec}}=-\sum_{t=1}^{l}\log p(z^{t}_{i}|X_{u},z^{1}_{i},z^{2}_{i},\cdots,z^{t-1}_{i}).$
In this work, we use a pretrained T5 (Raffel et al., 2020) as the language model. The parameters of T5 are also fine-tuned through back propagation to better adapt the language model for recommendation.
4.4. Item-Item Indexing
To align collaborative signals and item content information, we introduce an item-item indexing task which conducts the mapping from the content-based semantic space into the interaction-based collaborative space.
Model Inputs. The input sequence for item indexing contains the textual information of the item. Besides, we also introduce information of users who have interacted with this item, to further encode collaborative signals. Similar to the recommendation task, we augment the input an “item prompt” (denoted as $prompt_{i}$) at the beginning of the input. Therefore, the input of the indexing task is formulated as:
(7) $X_{i}=[prompt_{i},c_{i},\{uad_{u}|u\in\mathcal{U}_{i}^{+}\}].$
Language Models for Indexing. The indexing task is conducted through the same language model and codebook embeddings as the recommendation task. The generation probability for the indexing task is formulated similarly with Eq. (4) and Eq. (5) except that the model input is $X_{i}$ instead of $X_{u}$. We adopt the cross-entropy loss for parameter tuning. The loss for item indexing is defined as:
(8) $\mathcal{L}_{\text{index}}=-\sum_{t=1}^{l}\log p(z^{t}_{i}|X_{i},z^{1}_{i},z^{2}_{i},\cdots,z^{t-1}_{i}).$
Contrastive Optimization. To conduct better alignment between collaborative signals and content information, a contrastive loss is further introduced. The idea is that items with similar GIDs should also be similar in the content-based semantic space. To this end, for the item $i$, we randomly sample an item $i_{+}$ which has overlapped sub-sequence in GIDs as the positive sample, and another randomly sampled item $i_{-}$ without overlapped GID tokens as the negative sample. The contrastive loss is defined as:
(9) $\mathcal{L}_{\text{c}}=-\ln\sigma(\textbf{h}(X_{i})\cdot(\textbf{h}(X_{i_{+}})-\textbf{h}(X_{i_{-}}))),$
where $\mathbf{h}(\cdot)$ denotes the last hidden states of Encoder($\cdot$), and $\sigma$ denotes the sigmoid function. Such a contrastive loss helps the encoder to learn better item input representations.
4.5. Joint Optimization
Besides the above tasks, we further introduce a ranking loss to enhance the ranking ability of ColaRec. For a $(u,i)$ pair in the training dataset, we sample one item that the user $u$ has not interacted with, and for which the GID has no overlap with the positive item $i$, as the negative sample $i_{-}$. The BPR loss (Rendle et al., 2012) is utilized to optimize the ranking, which is formulated as:
(10) $\mathcal{L}_{\text{BPR}}=-\ln\sigma(\mathbf{h}(X_{u})\cdot(\mathbf{h}(X_{i})-\mathbf{h}(X_{i_{-}}))).$
where $\mathbf{h}(\cdot)$ denotes the last hidden states of Encoder($\cdot$). The above loss pushes together the positive $(u,i)$ pair in the semantic space and pushes away the negative $(u,i_{-})$ pair, thus helping the encoder to capture the ranking knowledge between items.
Finally, ColaRec is trained with the above-described tasks jointly:
(11) $\mathcal{L}=\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{index}}+\mathcal{L}_{\text{bpr}}+\alpha\mathcal{L}_{\text{c}},$
where $\alpha$ denotes the weight for the contrastive loss.
During the inference, recommendation lists are generated via beam search. To avoid the recommender from generating invalid GIDs that cannot be mapped to concrete items in the candidate set, we employ the constrained beam search (Cao et al., 2021) to limit the generative range of the current token based on the prefix tokens.
Table 1. Statistics of three public datasets after preprocessing.
Datasets #Users #Items #Interactions
Beauty 22,363 12,101 198,502
Sports 35,598 18,357 296,337
Phone 27,879 10,429 194,439
Table 2. Performance comparison on three public datasets. The best and the second-best scores are marked in bold and underlined fonts, respectively. * denotes the paired t-test with significance p-value ¡ 0.1.
CF-based Methods Generative Methods
Datasets Metric NeuMF LightGCN SimpleX NCL MutiVAE DiffRec DSI-R DSI-S TIGER Ours
Beauty Recall@5 0.0447 0.0649 0.0551 0.0650 0.0530 0.0524 0.0128 0.0451 0.0519 0.0667*
Recall@10 0.0653 0.0952 0.0831 0.0940 0.0776 0.0741 0.0228 0.0705 0.0799 0.0993*
Recall@20 0.0889 0.1314 0.1193 0.1320 0.1093 0.1016 0.0360 0.1018 0.1154 0.1371*
NDCG@5 0.0315 0.0450 0.0377 0.0452 0.0362 0.0378 0.0084 0.0305 0.0350 0.0449
NDCG@10 0.0383 0.0549 0.0469 0.0547 0.0443 0.0450 0.0117 0.0385 0.0443 0.0556*
NDCG@20 0.0445 0.0643 0.0563 0.0646 0.0526 0.0521 0.0151 0.0470 0.0534 0.0654*
Sports Recall@5 0.0206 0.0418 0.0355 0.0427 0.0314 0.0273 0.0117 0.0320 0.0374 0.0442*
Recall@10 0.0321 0.0623 0.0557 0.0631 0.0476 0.0403 0.0178 0.0497 0.0572 0.0660*
Recall@20 0.0471 0.0901 0.0836 0.0908 0.0713 0.0569 0.0284 0.0766 0.0881 0.0964*
NDCG@5 0.0140 0.0288 0.0240 0.0294 0.0208 0.0193 0.0079 0.0225 0.0249 0.0294
NDCG@10 0.0177 0.0355 0.0306 0.0361 0.0261 0.0235 0.0099 0.0284 0.0313 0.0364
NDCG@20 0.0215 0.0426 0.0377 0.0431 0.0321 0.0278 0.0126 0.0350 0.0392 0.0442*
Phone Recall@5 0.0410 0.0713 0.0643 0.0717 0.0569 0.0470 0.0187 0.0412 0.0601 0.0745*
Recall@10 0.0603 0.1052 0.0976 0.1043 0.0855 0.0668 0.0341 0.0625 0.0895 0.1121*
Recall@20 0.0871 0.1487 0.1420 0.1481 0.1233 0.0928 0.0564 0.0966 0.1299 0.1587*
NDCG@5 0.0282 0.0481 0.0423 0.0488 0.0378 0.0315 0.0121 0.0282 0.0403 0.0490
NDCG@10 0.0344 0.0590 0.0530 0.0593 0.0470 0.0379 0.0170 0.0347 0.0498 0.0611*
NDCG@20 0.0412 0.0700 0.0643 0.0704 0.0566 0.0445 0.0225 0.0431 0.0600 0.0729*
5. Experiment Settings
In this section, we describe experiment settings to evaluate the proposed ColaRec. We aim to answer the following research questions:
  1. RQ1
    How does the proposed ColaRec perform compared with existing recommendation methods?
  2. RQ2
    How does the joint training of multiple tasks affect the performance of ColaRec?
  3. RQ3
    How does the design of GIDs affect the recommendation performance?
5.1. Datasets
We use three real-world public datasets from Amazon Product Reviews [^3] https://jmcauley.ucsd.edu/data/amazon/, one of the most widely used recommendation benchmarks, to evaluate the performance of ColaRec. In particular, the experiments are conducted on three subcategories, including “Beauty”, “Sports and Outdoors”, and “Cell Phones and Accessories”. Users and items which have less than five interactions are filtered out. Table 1 shows the statistics of all three datasets after preprocessing. As for content information, we select “title”, “brand” and “categories” from the item metadata as the textual content description of items.
5.2. Evaluation Protocols
We adopt cross-validation to evaluate the performance of recommnders. In this paper we target on general recommendation other than sequential recommendation. To this end, we randomly split each user’s historical interactions into the training/validation/test set with the ratio of 8:1:1. We employ two widely used metrics, recall@$n$ and normalized discount cumulative gain (NDCG@$n$), to evaluate the model performance. Recall evaluates how many ground-truth items occur in the recommended list, while NDCG further focuses on their rankings in the list. Note that in this paper, the candidate item set is the whole item set, other than a small subset with selected items. Each experiment is conducted three times and the average score is reported.
5.3. Baselines
We compare ColaRec with several representative related baselines, including both conventional CF-based methods and generative methods. CF-based baselines include NeuMF (He et al., 2017), LightGCN (He et al., 2020), SimpleX (Mao et al., 2021), and NCL (Lin et al., 2022). Generative baselines include MultiVAE (Liang et al., 2018), DiffRec (Wang et al., 2023b), DSI (Tay et al., 2022), and TIGER (Rajput et al., 2023).
Table 3. Performance comparison of long-tail users on three public datasets. The best and the second-best scores are marked in bold and underlined fonts, respectively. ** denotes that the improvements are significant with p-value ¡ 0.05.
CF-based Methods Generative Methods
Datasets Metric NeuMF LightGCN SimpleX NCL MutiVAE DiffRec DSI-R DSI-S TIGER Ours
Beauty Recall@5 0.0416 0.0636 0.0555 0.0639 0.0510 0.0464 0.0131 0.0415 0.0519 0.0660**
Recall@10 0.0604 0.0922 0.0825 0.0907 0.0742 0.0662 0.0228 0.0653 0.0799 0.0975**
Recall@20 0.0817 0.1253 0.1160 0.1264 0.1039 0.0917 0.0354 0.0940 0.1154 0.1327**
Sports Recall@5 0.0209 0.0433 0.0355 0.0440 0.0329 0.0267 0.0116 0.0307 0.0380 0.0456**
Recall@10 0.0317 0.0639 0.0562 0.0645 0.0495 0.0394 0.0170 0.0472 0.0581 0.0674**
Recall@20 0.0468 0.0904 0.0836 0.0908 0.0725 0.0553 0.0273 0.0728 0.0882 0.0976**
Phone Recall@5 0.0405 0.0723 0.0660 0.0727 0.0571 0.0451 0.0206 0.0404 0.0602 0.0756**
Recall@10 0.0590 0.1054 0.0986 0.1043 0.0861 0.0641 0.0371 0.0623 0.0898 0.1131**
Recall@20 0.0855 0.1482 0.1418 0.1473 0.1228 0.0899 0.0600 0.0939 0.1293 0.1590**
  • NeuMF (He et al., 2017) enhances MF with multiple hidden layers to learn non-linear patterns in user-item interactions.
  • LightGCN (He et al., 2020) simplifies GNN for CF and learns user and item representations via linear neighborhood aggregation.
  • SimpleX (Mao et al., 2021) is a simple CF model with a cosine-based contrastive loss and negative sampling.
  • NCL (Lin et al., 2022) improves LightGCN with contrastive learning.
  • MultiVAE (Liang et al., 2018) is an autoencoder-based method, which utilizes VAEs to model the interaction signals.
  • DiffRec (Wang et al., 2023b) is a new proposed recommendation model based on diffusion models. DiffRec learns the user-item interaction knowledge through a recontruction and denoising manner.
  • DSI (Tay et al., 2022) is a generative document retrieval method. To adapt DSI for recommendation, we formulate the input as GIDs of the user’s historical interacted items. We use two versions of DSI. DSI-R refers to the DSI model with a random string as the GID of the item. DSI-S is a DSI model which constructs the item GID with a hierarchical $K$-means algorithm based on the item textual content embeddings from a pre-trained BERT model.
  • TIGER (Rajput et al., 2023) is a generative recommendation method. Specifically, a pre-trained Sentence-T5 encoder is used to obtain embeddings of the item’s textual content. These embeddings are then quantized using an RQ-VAE to build GIDs. We do not introduce sequential orders to adapt TIGER for general recommendation.
We don’t introduce P5-based baselines  (Geng et al., 2022; Hua et al., 2023) since these methods require the model input prompt to include candidate items for the general recommendation task. Given the limited input length, these methods cannot perform item ranking among the whole item set.
5.4. Implementation Details
[…]
Table 4. Ablation study on three datasets. The best and the second-best scores are marked in bold and underlined fonts, respectively. R@{5,10} and N@{5,10} stand for Recall@{5,10} and NDCG@{5,10} respectively.
Beauty Sports Phone
R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
ColaRec 0.0667 0.0993 0.0449 0.0556 0.0442 0.0660 0.0294 0.0364 0.0745 0.1121 0.0490 0.0611
(1) w/o textual content 0.0527 0.0809 0.0346 0.0439 0.0364 0.0542 0.0239 0.0297 0.0636 0.0974 0.0426 0.0535
(2) w/o indexing 0.0637 0.0947 0.0428 0.0531 0.0422 0.0644 0.0278 0.0350 0.0728 0.1086 0.0487 0.0602
(3) w/o $\mathcal{L}_{bpr}$ 0.0612 0.0918 0.0412 0.0512 0.0424 0.0634 0.0282 0.0350 0.0719 0.1077 0.0486 0.0602
(4) w/o $\mathcal{L}_{c}$ 0.0657 0.0989 0.0434 0.0544 0.0422 0.0649 0.0279 0.0352 0.0731 0.1106 0.0485 0.0606
Table 5. Comparison between different types of GIDs. The best and the second-best scores are marked in bold and underlined fonts, respectively. R@{5,10} and N@{5,10} stand for Recall@{5,10} and NDCG@{5,10} respectively.
Beauty Sports Phone
R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
ColaRec 0.0667 0.0993 0.0449 0.0556 0.0442 0.0660 0.0294 0.0364 0.0745 0.1121 0.0490 0.0611
(1) $iad$ 0.0658 0.0983 0.0437 0.0544 0.0428 0.0657 0.0285 0.0359 0.0719 0.1074 0.0474 0.0589
(2) $Random$ 0.0600 0.0902 0.0401 0.0500 0.0411 0.0623 0.0272 0.0340 0.0667 0.1004 0.0443 0.0551
(3) $Content$ 0.0662 0.1008 0.0440 0.0554 0.0423 0.0647 0.0278 0.0350 0.0716 0.1080 0.0477 0.0593
5.5. Performance Comparison (RQ1)
To answer RQ1, we conduct a comparative analysis of the proposed ColaRec on both overall users and long-tail users.
5.5.1. Comparison on whole users
Table 2 shows the performance comparison on overall users. From these results, we make the following observations. Firstly, the proposed ColaRec achieves the best recommendation performance on all datasets except for NDCG@5 in Beauty, which achieves comparable scores with the best NCL baseline. In particular, ColaRec consistently outperforms previous CF-based and generative baselines in Recall@20, achieving a relative improvement of 3.87%, 6.17% and 6.72% on Beauty, Sports, and Phone, respectively. These results demonstrate the effectiveness of ColaRec and its generalization across different domains.
Secondly, compared with DSI-R and DSI-S, which directly adapt generative retrieval methods to the recommendation task, ColaRec achieves a notable 47.89%, 38.13%, and 44.70% relative improvement in terms of Recall@5 on three datasets respectively, demonstrating effectiveness of the proposed ColaRec. Such results also demonstrate that naively transferring the generative retrieval methods for recommendation cannot achieve satisfying results. Furthermore, our method consistently outperforms the state-of-the-art generative recommendation model TIGER, which uses a RQ-VAE module to construct GIDs. The reason may be that TIGER only considers item content information but the collaborative signals are overlooked without an explicit alignment process.
Lastly, existing generative methods (e.g., DiffRec and TIGER), while exhibiting promising results in some specific scenarios (Rajput et al., 2023; Wang et al., 2023b; Li et al., 2023a), still underperform the strong CF-based methods (e.g., NCL) in the general recommendation task. In contrast, ColaRec achieves competitive results compared to these CF methods on all datasets, demonstrating the potential of infusing content information for collaborative generation of recommender systems.
5.5.2. Comparison on long-tail users
We also conduct experiments to verify the recommendation performance of ColaRec on long-tail users with sparse interactions. In this experiment, the ratio between head users and long-tail users is set as 20%:80%. Table 3 reports the results of Recall on three datasets. Results of NDCG show similar trends and are omitted due to the reason of space. We can see that ColaRec significantly outperforms all baselines when generating recommendation for long-tail users. The reason is that ColaRec models both user-item interactions and item content information. Given the long-tail users with less interaction knowledge, the content information helps ColaRec to gain better performance.
To conclude, the proposed ColaRec is effective to yield better performance compared with existing baselines. This improvement is more significant on long-tail users.
5.6. Ablation Study (RQ2)
In this section, we conduct ablation studies to analyse the effectiveness of each component in ColaRec. We implement four ablative variants of ColaRec, including: (1) w/o textual contentdeletes all textual content information in the model input; (2) w/o indexingremoves the item-item indexing task; (3) w/o $\mathcal{L}_{bpr}$removes the ranking loss $\mathcal{L}_{\text{bpr}}$; and (4) w/o $\mathcal{L}_{c}$removes the contrastive loss $\mathcal{L}_{\text{c}}$.
The results on three datasets are shown in Table 4. These results clearly indicate that the removal of any component from our proposed method results in a noticeable decline in overall performance. From the results of variant (1), we see that removing textual content information leads to a significant reduction in performance. Specifically, a $26.57\%$, $21.43\%$, and $17.14\%$ drop in Recall@5 is observed across the three datasets. This suggests the importance of incorporating textual content information to enhance the models’ understanding of items. Furthermore, the variant (3), which is trained without $\mathcal{L}_{\text{bpr}}$, exhibits a notable performance decrease in comparison to ColaRec. This highlights the effectiveness of the pairwise ranking objective, which focuses on the relative prioritization of positive and negative items within the generative recommendation paradigm. Besides, the results of variant (2) and (4) verify the effectiveness of the proposed item-item indexing task and the contrastive objective $\mathcal{L}_{c}$. A consistent performance reduction is observed on the three datasets when either of the two techniques is removed. This suggests that aligning item content information with user-item collaborative signals not only facilitates mutual reinforcement but also enables the learning of more comprehensive and effective representations.
In conclusion, each component of ColaRec is essential to improve the recommendation performance.
(a) Beauty
(b) Sports
(c) Phone
Figure 4. Impact of the length of GIDs. The GID length $l$ is set from $1$ to $4$.
(a) Beauty
(b) Sports
(c) Phone
Figure 5. Impact of the number of clusters. The number of clusters $K$ varies from $32$ to $128$.
5.7. GID Investigation (RQ3)
The construction of GIDs plays a crucial role in generative recommendation as it defines the model’s search space for generation. To answer RQ3, we conduct two analytical experiments about GIDs: (1) an analysis of different GID types (e.g., the single token $iad$, the random assigned GID, and the content-based GID), and (2) an analysis of hyper-parameters when constructing GIDs (e.g., the length and $K$-means cluster numbers).
5.7.1. Effect of different GID types
To further evaluate our GID generation strategy, we conducted an ablation study comparing it with three techniques: iad-based GID, $Random$ GID, and $Content$ GID. Specifically, the iad-based GID assigns a unique single token to each item to represent it using the corresponding vector in the item embedding matrix. The Random GID assigns a random string to each item as the identifier, without considering any prior knowledge. Content GID represents the GID constructed from item content information. Unlike the collaborative GID used in ColaRec, the content-based GID employs a hierarchical $K$-means clustering algorithm (Tay et al., 2022) to cluster items based on their textual representation derived from a pretrained BERT. To make a fair comparison with ColaRec, the length and the codebook size of both $Random$ GIDs and $Content$ GIDs are identical to those in ColaRec.
Table 5 details our results on the three datasets, where ColaRec attains the best performance. This indicates that our GID construction method, based on collaborative signals, is highly effective for generative recommendation. The improvement over the Content GID highlights the effectiveness of collaborative signals in recommendation tasks. Furthermore, ColaRec’s superior performance compared to the iad method demonstrates the benefit of explicitly introducing item correlations into item GIDs. In addition, the Random method leads to the lowest performance as the random string could further introduce noisy signals in the learning process. These results illustrate the importance of constructing effective GIDs for generative recommenders.
5.7.2. Impact of GID Hyper-parameters
In this section, we investigate the impact of hyperparameters $l$ and $K$ in the GID construction process. Figure 4 shows the results of ColaRec with different GID lengths $l$, ranging from $1$ to $4$. We can see that as $l$ varies, the recommedation performance exhibits some fluctuations, with the best results typically observed when $l=3$. It is worth noting that the GID of ColaRec consists of two parts: (i) the prefix tokens $[z_{i}^{1},\dots,z_{i}^{l-1}]$ from hierarchy clustering; and (ii) the last token $z_{l}$, which is randomly assigned to ensure uniqueness of the GID. A longer GID tends to include a greater proportion of hierarchy clustering parts, potentially facilitating model learning. However, a too long GID means more autoregressive decoding steps in model generation, which increases search difficulty and inference latency. Therefore, balancing efficiency and effectiveness, we have selected $l=3$ as the default setting.
To assess how the number of clusters $K$ affects performance, we fixed the GID length $l$ as $l=3$ and varied $K$ with values $32$, $64$, $96$, and 128. As shown in Figure 5, a higher $K$ typically results in a slight decrease in overall performance. The decrease on Beauty is more notable. The reason is that a higher $K$ indicates a larger search space in the decoding process, and thus increase the generation difficulty. To this end, choosing a suitable clustering number according to number of items is important for generative recommendation.
6. Conclusions and future work
This paper proposes ColaRec, a novel framework to conduct content-based collaborative generation for recommender systems. As an end-to-end generative recommender, ColaRec effectively integrates both item content information and user-item collaborative signals within a unified framework through fine-tuning an encoder-decoder based language model. In addition, an auxiliary item indexing task and a contrastive loss are proposed to further align the model’s representations of item content and user-item collaborative signals. We have conducted extensive experiments on three real-world datasets and empirical results demonstrate the effectiveness of ColaRec.
In the future, we intend to investigate more methods to construct appropriate generative identifiers (GIDs) and adopt more effective approaches to better align content information and collaborative signals. We also plan to introduce larger language models together with larger volume of training data to generate better recommendation. Negative sampling for generative recommendation is one of our future works, either using GIDs to sample more informative negative samples or utilizing generative models to generate simulated negative instances. Besides, parameter efficient fine-tuning and model efficiency are also potential research directions.